OpenStudy (anonymous):

Show how ((cos^2(x)-sin^2(x))/sin(x))-2= sin^2(x)

OpenStudy (anonymous):

Yeah can someone show me the steps or what trig identities you used.

OpenStudy (anonymous):

\[\frac{\cos^2(x)-\sin^2(x)}{\sin(x)}-2=\sin^2(x)\]

OpenStudy (anonymous):

it's not you, latex is not rendering properly.

OpenStudy (anonymous):

You know you can edit your question right?

myininaya (myininaya):

I see the latex just fine

myininaya (myininaya):

\[\frac{\cos^2(x)-\sin^2(x)}{\sin(x)-2}=\sin^2(x) ?\]

OpenStudy (anonymous):

Aha, hard refresh works!

OpenStudy (anonymous):

I'm doing eigen values and eigen functions so the left expression has to equal the right.

myininaya (myininaya):

You said you wrote it wrong? Did I write it right?

OpenStudy (anonymous):

Sorry I deleted the wrong one already. The one I wrote is right. Anyways I should rephrase my questions. I want the left side to equal the right side. How can I change the left side in order to make the equation true so that the left hand side equals the right hand side?

myininaya (myininaya):

\[\frac{1-\sin^2(x)-\sin^2(x)}{\sin(x)}-2=\frac{1-2\sin^2(x)}{\sin(x)}-2=\frac{1-2\sin^2(x)-2\sin(x)}{\sin(x)}\] hmmm.... I don't think they are the same .... Did you see if you could find a counterexample?

myininaya (myininaya):

\[x=\frac{\pi}{2}\] \[\frac{\cos^2(x)-\sin^2(x)}{\sin(x)}-2=\sin^2(x)\] \[\frac{\cos^2(\frac{\pi}{2})-\sin^2(\frac{\pi}{2})}{\sin(\frac{\pi}{2})}=\frac{0-1}{1}=-1\] But the other side is 1 when x=pi/2 Therefore both sides are not the same

myininaya (myininaya):

-1-2 does not equal 1

myininaya (myininaya):

no questions?

OpenStudy (anonymous):

What do you suggest I do in order to get them equal?

myininaya (myininaya):

they are not equal

myininaya (myininaya):

this is not an identity I gave you a counterexample above try pluggin' in pi/2 for x we do not get the same thing on both sides unless you meant to type something else